OpenStudy (anonymous):

integrate cos^(1/2)sin^3x dx from 0 to pi/2

zepdrix (zepdrix):

\[\Large \int\limits_0^{\pi/2}\sqrt{\cos x}\;\sin^3x\;dx\]So this is what we've got to solve? Hmm

OpenStudy (anonymous):

Yes, so far I've pulled out one of the sinx then changed the \[\sin ^{2}x \to (1-\cos ^{2}x)\]

OpenStudy (anonymous):

then I guess you do a u sub and u=cosx and du would be -sinx

OpenStudy (anonymous):

but I'm not sure where to go from there

zepdrix (zepdrix):

-du=sinx? looks like you're on the right track

zepdrix (zepdrix):

\[\Large \int\limits_0^{\pi/2}\cos^{1/2}x\left(1-\cos^2x\right)(\sin x\;dx)\] \[\Large u=\cos x \qquad\to\qquad -du=\sin x\;dx\]

zepdrix (zepdrix):

If I wrote it like this, does it become a tad clearer?\[\Large -du=(\sin x\;dx)\]

zepdrix (zepdrix):

Maybe you can match it up with the problem that way hehe

OpenStudy (anonymous):

ahh, I see what you're saying

OpenStudy (anonymous):

then the negative on that would just come out from like a constant

zepdrix (zepdrix):

I would probably distribute the negative to each term in the (1-u^2) brackets. But yah, you can pull it out of the integral if you want.

OpenStudy (anonymous):

and the dx's would cancel. So you'd be left with \[-\int\limits_{0}^{\pi/2} u ^{1/2}(1-u ^{2}) du\]

zepdrix (zepdrix):

Looks good, let's just not forget that these limits of integration don't apply to our u variable.\[\Large -\int\limits\limits_{\color{red}{x=0}}^{\color{red}{\pi/2}} u ^{1/2}(1-u ^{2}) du\]

zepdrix (zepdrix):

writing x= is a good way to remind yourself of that.

OpenStudy (anonymous):

oh, okay. Yeah, I usually forget that :/

zepdrix (zepdrix):

So what's the next step? \c:/

OpenStudy (anonymous):

I would probably distribute the \[u ^{1/2}\] to the parenthesis then solve the equation like seperate integrals

zepdrix (zepdrix):

sounds good :O

OpenStudy (anonymous):

so now I have \[-(\frac{ 2 }{ 3 })u ^{3/2}-(\frac{ 1 }{ 3 })u ^{3}\]

OpenStudy (anonymous):

or did I forget something?

zepdrix (zepdrix):

Ah sorry I disappeared there :3 I would have probably distributed this negative into the brackets as well as distributing the u^1/2. I think you're less likely to make a mistake that way :o \[\Large -\int\limits\limits\limits_{\color{red}{x=0}}^{\color{red}{\pi/2}} u ^{1/2}(1-u ^{2}) du \qquad\to\qquad \Large \int\limits\limits_{\color{red}{x=0}}^{\color{red}{\pi/2}} u ^{5/2}-u^{1/2}\; du\]

zepdrix (zepdrix):

\[\Large u^{1/2}u^{2} \quad=\quad u^{5/2}\]

OpenStudy (anonymous):

apparently I can do calculus but can't add :/ lol

zepdrix (zepdrix):

hehe

OpenStudy (anonymous):

but did I have the rest correct? then I'd sub back for the u?

zepdrix (zepdrix):

AHHHH stupid openstudy freezing >:(

zepdrix (zepdrix):

One other small mistake, when you distribute that negative, you should have gotten a +on the second term.\[\Large -\frac{2}{3}u^{3/2}\color{red}{+}\frac{2}{7}u^{7/2}\]

OpenStudy (anonymous):

yeah :/

OpenStudy (anonymous):

I can't complain too much though. Don't know what I'd do without it

zepdrix (zepdrix):

From there, ya let's undo our substitution, then we can evaluate the result at our limits.

OpenStudy (anonymous):

ah, okay

OpenStudy (anonymous):

\[\left[ -\frac{ 2 }{ 3 } \right]cosx ^{3/4}+\frac{ 2 }{ 7 }\cos ^{7/2}\] evaluated at 1 to 0?

OpenStudy (anonymous):

\[\left[ -\frac{ 2 }{ 3 } \right]cosx ^{3/4}+\frac{ 2 }{ 7 }\cos ^{7/2}\] evaluated at 1 to 0?

zepdrix (zepdrix):

The first one is to the 3/2 power right? :o

OpenStudy (anonymous):

yeah, typo

zepdrix (zepdrix):

Evaluated from 0 to pi/2. \[\Large \frac{2}{7}\cos^{7/2}x-\frac{2}{3}\cos^{3/2}x\;|_0^{\pi/2}\] But yes, you're right, our cosines are going to be producing a bunch of 0's and 1's.

OpenStudy (anonymous):

don't you have to change the limits?

zepdrix (zepdrix):

If we were going to evaluate this while it was in u, then yes. \[\Large \frac{2}{7}\cos^{7/2}x-\frac{2}{3}\cos^{3/2}x\;|_{\color{green}{x=0}}^{\color{green}{\pi/2}}\] Or, \[\Large \frac{2}{7}u^{7/2}-\frac{2}{3}u^{3/2}\;|_{u=1}^{0}\]

zepdrix (zepdrix):

When you do a substitution, you have 2 options: 1. You can change the limits along with doing your substitution. 2. You can undo your substitution at the end, before you plug in the limits, in which case you don't need to change the limits.

OpenStudy (anonymous):

oh, nice. That's part of what has been confusing me. So after I change it back from the u I can just plug in the original limits?

zepdrix (zepdrix):

ya c:

zepdrix (zepdrix):

You have the option to NOT CHANGE IT BACK TO U, that's the case where you would have to figure out new limits. Both methods work just fine :o

OpenStudy (anonymous):

awesome, this problem cleared up a lot of confusion. I got 8/21 for the final answer

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

awesome! Thank you so much!